1)The perimeter of a rectangle field is 760m. If the length is 20m longer than the breadth, find the dimension of both the length and the breadth
2) A father is 15 year older than his son, the sum of their age is 45years.find the age

Answer by Cromlix(4381)

Hi there,
1)The perimeter of a rectangle field is 760m. If the length is 20m longer than the breadth, find the dimension of both the length and the breadth.
Make the breadth = x
length = 20 + x
Perimeter = 2 x breadth + 2 x length
760 = 2(x) + 2(20 + x)
Multiply out
760 = 2x + 40 + 2x
Collect like terms:
2x + 2x = 760 - 40
4x = 720
x = 180
Breadth = 180 m
Length = 200 m
...................
2) A father is 15 year older than his son, the sum of their age is 45years.find the age
Father's age = 15 + Son's age.
F = 15 + S
the sum of their age is 45years.
(15 + S) + S = 45
Remove brackets
15 + S + S = 45
Collect like terms
S + S = 45 - 15
2S = 30
S = 15
Son's age is 15 years old
Father's age is 30 years old
Hope this helps :-)